# Specific heat of Gd3+ and Eu2+-based magnetic compounds

### v4

@@ Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, corresponding to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (determined by the magnitude of the exchange interactions), we find that, to a high degree of accuracy, the specific heat is governed by two primary parameters: the effective number of neighbors $z$, which dictates the extent of spatial and quantum fluctuations, and the axial anisotropy KK. The universality of $z$ (its ability to describe specific heat across diverse lattices) holds robustly for systems where exchange interactions do not strongly increase with distance and in the absence of frustration. Otherwise, deviations from universality emerge. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds, achieving a remarkable agreement. The present approach enables the extraction of magnetic interaction parameters not accessible through mean-field theory, offering a powerful tool for interpreting specific heat data in 4f7 systems.
@@ I Introduction
Recent research on Eu2+-based compounds shows similar shortcomings of the mean-field approach [15, 16]. The fact that Eu2+ has a much larger atomic volume suggests it may participate in the formation of distinct crystalline structures. This indicates that the observed deviations are an intrinsic characteristic of the magnetic interactions. They are not dependent on crystalline symmetries or the sign of the interactions. Therefore, for a quantitative explanation of the observed $C(T)$, it is necessary to include spatial and quantum fluctuations and go beyond mean field. The effect of these fluctuations was calculated recently using cluster mean-field theory [17].
@@ II.1 The effective number of neighbors
We can draw an analogy to a binary random distribution (which would correspond to spin 1/2). It is known that after $N$ attempts in a binary random distribution $x$ with probabilities $p$ and 1p1-p for the values $x=1$ and x=1x=-1, respectively, the mean value (proportional to the magnetic moment) is $\langle x\rangle=(2p-1)N$, and the standard deviation is $\sqrt{\langle x^{2}\rangle-\langle x\rangle^{2}}=2\sqrt{Np(1-p)}$. Thus, the ratio between the standard deviation and the mean value decreases as $1/\sqrt{N}$. Similarly, in our quantum Heisenberg model, one expects that as the number of neighbors increases, the specific heat is more similar to the mean-field result, in which spatial and quantum fluctuations are neglected and the effective magnetic field that each ion feels is replaced by its mean value.
Generalizing the above results to the case of a distribution with intensities $\pm|J_{\delta}|$ for $z_{\delta}$ neighbors, we obtain after some algebra (see Appendix):
@@
Here, $z_{\delta}$ represents the number of neighbors of type $\delta$ (with exchange interaction $J_{\delta}$), and the summation goes over all types of neighbors. The quantity $z$ quantifies the effective number of neighbors, taking into account the strength ($|J_{\delta}|$) and number ($z_{\delta}$) of each type of interaction. Remarkably, for unfrustrated systems, the observed specific heat depends mainly on this effective number of neighbors $z$, largely independent of the specific lattice structure. $z$ quantifies the relative strength of spatial fluctuations. For larger $z$ (higher coordination number), the results approach the mean-field ones, where spatial and quantum fluctuations are suppressed. Conversely, for lower $z$ (fewer neighbors), fluctuations become more important, leading to a smaller specific heat below the critical temperature $T_{C}$ and a larger specific heat above $T_{C}$. Notably even for $z$ as large as $26$ there is a large departure of the quantum specific heat respect to the mean field result.
@@ II.2 Axial anisotropy in 4f7 ions
Nevertheless, due to spin-orbit coupling, the ground-state multiplet acquires a small admixture of the $^{6}P_{7/2}$ state with $L=1$. An explicit calculation for Gd+3 in Appendix B of Ref. 18 estimates that the excited multiplet contributes approximately $3\%$ to the ground state. This admixture gives rise to an $L=1$ component in the ground state, which enables its splitting in the presence of crystal fields. This occurs if the symmetry at the rare-earth site is not high. Specifically, if there is only a single axis of higher order, the anisotropy term $H_{A}=C[3L_{z}^{2}-L(L+1)]$ becomes relevant.
@@ III Methods
We have calculated the specific heat of a large number of magnetic structures. The approach used a two-point numerical differentiation to compute the derivative of the internal energy as a function of temperature. The internal energy was calculated using quantum Monte Carlo (QMC) simulations with the default “minbounce” algorithm (implemented in the “dirloop_sse” application in the ALPS libraries [19, 20]). QMC allows us to handle large system sizes and complex lattices while properly accounting for spatial and quantum fluctuations that mean-field approaches miss. Simulations were performed on systems of up to $8\times 30^{3}=216000$ magnetic moments. A typical simulation required approximately 10000 CPU hours per lattice on a standard workstation.
To generate the lattice configurations for the ALPS simulations, we employed the “SUNNY” package [21] in conjunction with crystallographic information files (CIF) obtained from public repositories. SUNNY reads the CIF file and outputs the list of neighbour vectors up to a chosen cutoff. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, MgAgAs (for GdPbBi), Th(CrSi)2 (for GdCu2Ge2), ErNiSi3 (for GdNiSi3), Mg2MnGa3 (for Eu2Pd3Sn) and MgAl2Cu (for EuPdSn2) were retrieved from the Materials Project [22] database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database [23] using data from Ref. 24. Lattice parameters for GdCu2Ge2 were taken from Ref. [25], for GdPbBi from Ref. [26], for GdNiSi3 from Ref. [27], for Eu2Pd3Sn from Ref. [28], and for EuPdSn2 from Ref. [29]. In each case, the magnetic structure was constructed by assigning a spin moment $S=7/2$ to the respective atoms (Gd or Eu). For all lattices, except HCP, we used a rectangular cuboid supercell geometry as the conventional unit cell, resulting in varying numbers of sites within the supercell, with the diamond lattice requiring the largest (8 sites). Simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, 2020, and $30$. For the critical temperature, results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
@@
Our theoretical results for the specific heat of all lattices used here are available in Ref. 31.
@@ IV Role of the effective number of neighbors
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\mathrm{MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section V.1). But it is lower than $T_{C}^{\mathrm{MF}}$, particularly for low $z$ due to the effect of spatial and quantum fluctuations, which is larger for low $z$ (a similar behavior was reported for $S=1/2$ in Ref. [32]).
@@
In Fig. 2 we show the results for the specific heat $C$ as a function of temperature for a large number of ferromagnetic structures with different numbers of magnetic neighbors $z$ (the number of nearest neighbors up to third order is shown in Table 1 and its critical temperatures are shown in Table 2). Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, the specific heat for a honeycomb layered system with $z=4$ (three nearest neighbors within the plane of the honeycomb layer and one above or below the layer in alternating sites, LHc(4)) shows the smallest specific heat below $T_{C}$ and the largest one above $T_{C}$. Note that the diamond structure, which is distinct from the previous one but also has $z=4$, has a very similar specific heat. The low-temperature shoulder is a Schottky-like effect due to the internal field [33, 14]. The lower panel inset shows that finite size effects are only appreciable near $T_{C}$ as exemplified in the case of the HCP(12) lattice. For this reason, we show specific heat curves from QMC corresponding to $N=20$.
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the BCC structure with 6 NN and 8 next-nearest neighbors (NNN), the same trend is observed in general. However, the specific heat of the BCC lattice with $z=8$ (first neighbors) lies very close to the cubic FCC with $z=12$ and above $z=10$. The FCC lattice with $z=12$ has a $C$ quite close to the HCP with $z=12$ (first and second neighbors). The diamond(16) and EuPdSn2(16) lattices when 16 neighbors are considered also have very similar specific heat. Other exception correspond to the FCC(18) case composed of twelve NN and six NNN that has $C$ quite close to the previous mentioned cases of diamond(16) and EuPdSn2(16). Nevertheless, the specific heat is very similar for different structures with the same $z$. Even if the mean-field theory predicts identical behavior for all coordination numbers, without accounting for lattice geometry, this universality seen with QMC across different lattice types with the same $z$ is remarkable.
@@ V.1 Mean-field Approximation
The mean-field approximation assumes that all sites are equivalent (long-range magnetic order), which may not hold near the ordering temperature. We assume that there is no external magnetic field. For simplicity, we restrict the analysis to the easy axis (uniaxial anisotropy or anisotropy along a single axis) case $K<0$, for which the spins point in the direction of the anisotropy axis $z$. In the next subsection, both signs of $K$ and the effects of spatial and quantum fluctuations will be considered.
@@ V.2 Beyond Mean-field
An anisotropy-induced energy gap arises from the splitting of spin states due to the axial anisotropy term $H_{A}=K[3S_{z}^{2}-S(S+1)]$. For negative $K$, this term favors alignment along the $z$-axis, creating a gap between low-energy $|S_{z}|$ states with large $|S_{z}|$ values. This restricts spin fluctuations above $T_{C}$, reducing entropy and specific heat. For positive $K$, the anisotropy favors in-plane alignment, leading to a inverse distribution of low-energy states but with similar effects on the specific heat. The QMC simulations explicitly capture these quantum and spatial fluctuations, which mean-field theory cannot account for.
@@ VI Comparison of Different Structures with the Same Effective Number of Neighbors
Our central finding is that the specific heat of magnetic systems is primarily determined by the effective number of neighbors $z$, defined by Eq. (4) rather than the specific lattice structure or interaction pattern. The parameter $z$ quantifies the relative strength of spatial and quantum fluctuations. Physically, $z$ represents the effective number of neighbors that meaningfully contribute to the exchange field experienced by each spin, accounting for both geometric coordination and relative interaction strengths. In Section IV, we analyzed the impact of varying $z$ on the specific heat, while keeping $J_{\delta}=J_{1}$. Here, we extend this analysis by investigating how changes in $J_{\delta}$ affect the specific heat, while maintaining the effective number of neighbors $z$ constant [see Eq. (4)].
@@
These results demonstrate that the effective number of neighbors $z$ encapsulates the essential physics of spatial and quantum fluctuations for $0\geq J_{2}\geq J_{1}$ ($|J_{2}|\leq|J_{1}|$), and it provides a good approximation for $|J_{2}|>|J_{1}|$. This enables a universal description of specific heat across diverse magnetic structures, independent of their lattice geometry or interaction patterns.
@@ VI.1 Getting information beyond mean field
This observation provides a powerful tool for experimentalists. By comparing experimental specific heat data to our reference curves (available in Ref. 31), one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This is a significant advance over mean-field analysis, which cannot distinguish between different interaction patterns that yield the same critical temperature.

### v5

@@ Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, corresponding to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (determined by the magnitude of the exchange interactions), we find that, to a high degree of accuracy, the specific heat is governed by two primary parameters: the effective number of neighbors $z$, which governs the extent of thermal and quantum fluctuations, and the axial anisotropy KK. The universality of $z$ (its ability to describe specific heat across diverse lattices) holds robustly for systems where exchange interactions do not strongly increase with distance and in the absence of frustration. Otherwise, deviations from universality emerge. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds, achieving a remarkable agreement. The present approach enables the extraction of magnetic interaction parameters not accessible through mean-field theory, offering a powerful tool for interpreting specific heat data in 4f7 systems.
@@ I Introduction
Recent research on Eu2+-based compounds shows similar shortcomings of the mean-field approach [15, 16]. The fact that Eu2+ has a much larger atomic volume suggests it may participate in the formation of distinct crystalline structures. This indicates that the observed deviations are an intrinsic characteristic of the magnetic interactions. They are not dependent on crystalline symmetries or the sign of the interactions. Therefore, for a quantitative explanation of the observed $C(T)$, it is necessary to include thermal and quantum fluctuations and go beyond mean field. The effect of these fluctuations was calculated recently using cluster mean-field theory [17].
@@
In Appendix A we discuss the relation of $z$ with the quantum fluctuations. Finally, quantitative measures of the differences between specific-heat curves are presented in Appendices B, C, D, and E.
@@ II.1 The effective number of neighbors
We can draw an analogy to a binary random distribution (which would correspond to spin 1/2). It is known that after $N$ attempts in a binary random distribution $x$ with probabilities $p$ and 1p1-p for the values $x=1$ and x=1x=-1, respectively, the mean value (proportional to the magnetic moment) is $\langle x\rangle=(2p-1)N$, and the standard deviation is $\sqrt{\langle x^{2}\rangle-\langle x\rangle^{2}}=2\sqrt{Np(1-p)}$. Thus, the ratio between the standard deviation and the mean value decreases as $1/\sqrt{N}$. Similarly, in our quantum Heisenberg model, one expects that as the number of neighbors increases, the specific heat is more similar to the mean-field result, in which thermal and quantum fluctuations are neglected and the effective magnetic field that each ion feels is replaced by its mean value.
Generalizing the above results to the case of a distribution with intensities $\pm|J_{\delta}|$ for $z_{\delta}$ neighbors, we obtain after some algebra (see Appendix A):
@@
Here, $z_{\delta}$ represents the number of neighbors of type $\delta$ (with exchange interaction $J_{\delta}$), and the summation goes over all types of neighbors. The quantity $z$ quantifies the effective number of neighbors, taking into account the strength ($|J_{\delta}|$) and number ($z_{\delta}$) of each type of interaction. Remarkably, for unfrustrated systems, the observed specific heat depends mainly on this effective number of neighbors $z$, largely independent of the specific lattice structure. $z$ quantifies the relative strength of spatial fluctuations. For larger $z$ (higher coordination number), the results approach the mean-field ones, where thermal and quantum fluctuations are suppressed. Conversely, for lower $z$ (fewer neighbors), fluctuations become more important, leading to a smaller specific heat below the critical temperature $T_{C}$ and a larger specific heat above $T_{C}$. Notably even for $z$ as large as $26$ there is a large departure of the quantum specific heat respect to the mean field result.
We note that $z$ is distinct from the effective coordination number $\hat{z}$ commonly used in inorganic and organic chemistry [18], where contributions from neighboring atoms are summed with weights decaying exponentially with distance. In contrast, our definition of $z$ assigns weights determined by the exchange constants $J_{\delta}$, which may be enhanced at longer distances and are not bound by geometric decay laws. Thus, $z$ captures the dynamical influence of exchange interactions on spin fluctuations, rather than the static geometry of atomic arrangements.
@@ II.2 Axial anisotropy in 4f7 ions
Nevertheless, due to spin-orbit coupling, the ground-state multiplet acquires a small admixture of the $^{6}P_{7/2}$ state with $L=1$. An explicit calculation for Gd+3 in Appendix B of Ref. 19 estimates that the excited multiplet contributes approximately $3\%$ to the ground state. This admixture gives rise to an $L=1$ component in the ground state, which enables its splitting in the presence of crystal fields. This occurs if the symmetry at the rare-earth site is not high. Specifically, if there is only a single axis of higher order, the anisotropy term $H_{A}=C[3L_{z}^{2}-L(L+1)]$ becomes relevant.
@@ III Methods
We have calculated the specific heat of a large number of magnetic structures. The approach used a two-point numerical differentiation to compute the derivative of the internal energy as a function of temperature. The internal energy was calculated using quantum Monte Carlo (QMC) simulations with the default “minbounce” algorithm (implemented in the “dirloop_sse” application in the ALPS libraries [20, 21]). QMC allows us to handle large system sizes and complex lattices while properly accounting for thermal and quantum fluctuations that mean-field approaches miss. Simulations were performed on systems of up to $8\times 30^{3}=216000$ magnetic moments. A typical simulation required approximately 10000 CPU hours per lattice on a standard workstation. Classical Monte Carlo simulations fail to capture the low-temperature physics of quantum spin systems. By neglecting spin quantization and the Bose-Einstein statistics of magnons, they yield two hallmark errors: a finite residual entropy at $T=0$ (violating the third law) and a divergent $C/T$ as $T\to 0$. This artifact, documented in Ref. [22], persists up to $\sim 0.75\,T_{C}$ and obscures the magnetic signal in the temperature range $0.25<T/T_{C}<1.25$ critical for comparison with experiment. Quantum Monte Carlo, by contrast, correctly enforces spin quantization and recovers $C\to 0$ as $T\to 0$, enabling a faithful representation of the full specific heat profile across all temperatures.
To generate the lattice configurations for the ALPS simulations, we employed the “SUNNY” package [23] in conjunction with crystallographic information files (CIF) obtained from public repositories. SUNNY reads the CIF file and outputs the list of neighbour vectors up to a chosen cutoff. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, MgAgAs (for GdPbBi), Th(CrSi)2 (for GdCu2Ge2), ErNiSi3 (for GdNiSi3), Mg2MnGa3 (for Eu2Pd3Sn) and MgAl2Cu (for EuPdSn2) were retrieved from the Materials Project [24] database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database [25] using data from Ref. 26. Lattice parameters for GdCu2Ge2 were taken from Ref. [27], for GdPbBi from Ref. [28], for GdNiSi3 from Ref. [29], for Eu2Pd3Sn from Ref. [30], and for EuPdSn2 from Ref. [31]. In each case, the magnetic structure was constructed by assigning a spin moment $S=7/2$ to the respective atoms (Gd or Eu). For all lattices, except HCP, we used a rectangular cuboid supercell geometry as the conventional unit cell, resulting in varying numbers of sites within the supercell, with the diamond lattice requiring the largest (8 sites). Simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, 2020, and $30$. For the critical temperature, results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
@@
Our theoretical results for the specific heat of all lattices used here are available in Ref. 33.
@@ IV Role of the effective number of neighbors
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\mathrm{MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section V.1). But it is lower than $T_{C}^{\mathrm{MF}}$, particularly for low $z$ due to the effect of thermal and quantum fluctuations, which is larger for low $z$ (a similar behavior was reported for $S=1/2$ in Ref. [34]).
@@
In Fig. 2 we show the results for the specific heat $C$ as a function of temperature for a large number of ferromagnetic structures with different numbers of magnetic neighbors $z$ (the number of nearest neighbors up to third order is shown in Table 1 and its critical temperatures are shown in Table 2). Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, the specific heat for a honeycomb layered system with $z=4$ (three nearest neighbors within the plane of the honeycomb layer and one above or below the layer in alternating sites, LHc(4)) shows the smallest specific heat below $T_{C}$ and the largest one above $T_{C}$. This behavior arises from short-range correlations, which are absent in mean-field theory and persist for $T>T_{C}$, thereby contributing to the specific heat. The importance of these correlations increases as $z$ decreases. Since the total entropy per site is fixed at $\ln(7)$ (for $S=7/2$), this redistribution of entropy implies that $C/T$ decreases with decreasing $z$ for $T<T_{C}$. High-temperature series expansions [34] confirm that, well above $T_{C}$, the specific heat of three-dimensional Heisenberg magnets should exhibit a $1/T^{3}$ tail, consistent with our observed behavior. Note that the diamond structure, which is distinct from the previous one but also has $z=4$, has a very similar specific heat. The low-temperature shoulder is a Schottky-like effect due to the internal field [35, 14]. The lower panel inset shows that finite size effects are only appreciable near $T_{C}$ as exemplified in the case of the HCP(12) lattice. For this reason, we show specific heat curves from QMC corresponding to $N=20$.
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the BCC structure with 6 NN and 8 next-nearest neighbors (NNN), the same trend is observed in general. However, the specific heat of the BCC lattice with $z=8$ (first neighbors) lies very close to the cubic FCC with $z=12$ and above $z=10$. The FCC lattice with $z=12$ has a $C$ quite close to the HCP with $z=12$ (first and second neighbors).
In Fig.  9 of Appendix C we quantify the differences between the specific heat curves using the distance measure defined in Appendix B.
The diamond(16) and EuPdSn2(16) lattices when 16 neighbors are considered also have very similar specific heat. Other exception correspond to the FCC(18) case composed of twelve NN and six NNN that has $C$ quite close to the previous mentioned cases of diamond(16) and EuPdSn2(16). Nevertheless, the specific heat is very similar for different structures with the same $z$. Even if the mean-field theory predicts identical behavior for all coordination numbers, without accounting for lattice geometry, this universality seen with QMC across different lattice types with the same $z$ is remarkable.
@@ V.1 Mean-field Approximation
The mean-field approximation assumes that all sites are equivalent (long-range magnetic order), which may not hold near the ordering temperature. We assume that there is no external magnetic field. For simplicity, we restrict the analysis to the easy axis (uniaxial anisotropy or anisotropy along a single axis) case $K<0$, for which the spins point in the direction of the anisotropy axis $z$. In the next subsection, both signs of $K$ and the effects of thermal and quantum fluctuations will be considered.
@@ V.2 Beyond Mean-field
An anisotropy-induced energy gap arises from the splitting of spin states due to the axial anisotropy term $H_{A}=K[3S_{z}^{2}-S(S+1)]$. For negative $K$, this term favors alignment along the $z$-axis, creating a gap between low-energy $|S_{z}|$ states with large $|S_{z}|$ values. This restricts spin fluctuations above $T_{C}$, reducing entropy and specific heat. For positive $K$, the anisotropy favors in-plane alignment, leading to a inverse distribution of low-energy states but with similar effects on the specific heat. The QMC simulations explicitly capture these quantum and thermal fluctuations, which mean-field theory cannot account for.
@@ VI Comparison of Different Structures with the Same Effective Number of Neighbors
Our central finding is that the specific heat of magnetic systems is primarily determined by the effective number of neighbors $z$, defined by Eq. (4) rather than the specific lattice structure or interaction pattern. The parameter $z$ quantifies the relative strength of thermal and quantum fluctuations. Physically, $z$ represents the effective number of neighbors that meaningfully contribute to the exchange field experienced by each spin, accounting for both geometric coordination and relative interaction strengths. In Section IV, we analyzed the impact of varying $z$ on the specific heat, while keeping $J_{\delta}=J_{1}$. Here, we extend this analysis by investigating how changes in $J_{\delta}$ affect the specific heat, while maintaining the effective number of neighbors $z$ constant [see Eq. (4)].
@@
Quantitative measures of the difference in specific heats are provided in Tables 5 and 6 of Appendix D
These results demonstrate that the effective number of neighbors $z$ encapsulates the essential physics of thermal and quantum fluctuations for $0\geq J_{2}\geq J_{1}$ ($|J_{2}|\leq|J_{1}|$), and it provides a good approximation for $|J_{2}|>|J_{1}|$. This enables a universal description of specific heat across diverse magnetic structures, independent of their lattice geometry or interaction patterns.
@@ VI.1 Getting information beyond mean field
This observation provides a powerful tool for experimentalists. By comparing experimental specific heat data to our reference curves (available in Ref. 33), one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This is a significant advance over mean-field analysis, which cannot distinguish between different interaction patterns that yield the same critical temperature.
@@
For quantitative comparisons between specific heat curves across lattices, including distances under uniform and non-uniform couplings, as well as fits to experimental data, see Appendices B, C, D, and E.